OpenStudy (anonymous):

Solve for x: −2|x − 3| = −12 A: x = −3, x = 4.5 B: x = 3, x = 9 C: x = −3, x = 9 D: No solutions

OpenStudy (jamesj):

Step one: divide both sides by -2, what do you get?

OpenStudy (anonymous):

X-3= 0.16666666666666667

OpenStudy (anonymous):

@JamesJ

OpenStudy (jamesj):

-12/-2 = 6 Hence |x - 3| = 6

OpenStudy (anonymous):

hahaa but you said -2/-12 so thats what I did lol

OpenStudy (jamesj):

I said divide both sides by -2. Anyway, what's the next step?

OpenStudy (anonymous):

w/e hahah... um, add 3 on both sides?

OpenStudy (jamesj):

|x-3| = 6 are numbers x which are a distance 6 from 3. What numbers satisfy that relationship?

OpenStudy (anonymous):

9?

OpenStudy (jamesj):

9 is one number 6 from 3. Yes. Any others?

OpenStudy (anonymous):

not sure ahaha...

OpenStudy (jamesj):

Draw a number line. Mark out 0 and 3. Ask yourself again: what numbers are a distance 6 from 3? there are two such numbers.

OpenStudy (anonymous):

4 and 5?

OpenStudy (jamesj):

4 is a distance of 1 from 3. 5 is a distance of 2 from 3.

OpenStudy (anonymous):

just help me out man hahaha Im not sure what you mean?

OpenStudy (jamesj):

6 is a distance of 3 from 3, yes? 7 is at a distance of 4 from 3, yes? 2 is a distance 1 from 3, yes? 1 is a distance of 2 from 3, yes?

OpenStudy (anonymous):

yeah

OpenStudy (jamesj):

|x-3| measures the distance of x from 3. For example, if x = 9, |x-3| = |9-3| = |6| = 6

OpenStudy (jamesj):

What other number, x has the property that |x-3| = 6 ?

OpenStudy (anonymous):

umm....

OpenStudy (jamesj):

If |x-3| = 6 then either x - 3 = 6 or x - 3 = -6 hence x = ..... or .....

OpenStudy (anonymous):

9 and

OpenStudy (anonymous):

damn!! Im not even sure!

OpenStudy (jamesj):

0 If |x-3| = 6 then either x - 3 = 6 or x - 3 = -6 Now, x - 3 = 6 ==> x = 6 + 3 = 9 if x - 3 = -6 ==> x = -6 + 3 = ....

OpenStudy (anonymous):

-3

OpenStudy (jamesj):

Yes -3 and +3 are a distance of 6 apart.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so is the answer c?

OpenStudy (jamesj):

Looks like it

OpenStudy (anonymous):

okay thanks.. you want to help me on some more?

OpenStudy (jamesj):

No, sorry. I have some of my own work to do now.

OpenStudy (anonymous):

damn.. okay